OpenStudy (sette):

Check My Answer ? "Solve for s" -6>=(-26s+27)/6>-7 Write your answer as a compound inequality with integers, proper fractions, and improper fractions in simplest form. I put 63/26=

OpenStudy (amistre64):

*6, -27, /-26, flip signs

OpenStudy (amistre64):

looks right at a cursory glance

OpenStudy (sette):

so my answer is right ?

OpenStudy (anonymous):

From Mathematica:\[\frac{63}{26}\leq s<\frac{69}{26} \]Yes.